OpenStudy (anonymous):

The probability of two heads on a toss of two coins is .25. Is this theoretical or empirical probability?

OpenStudy (anonymous):

What do you think so far?

OpenStudy (anonymous):

empirical because it is based on experimental evidence? @Tom_Boy_Rebel

OpenStudy (dan815):

its theoretical

OpenStudy (dan815):

we start with the hypothesis that a coin is indeed random, that there is 1/2 chance of heads and 1/2 chance of tails

OpenStudy (dan815):

so when toss twice there is a 1/2 * 1/2 chance of heads 1/2*1/2=1/4 = 0.25 However, this not could really be true, perhaps not all coins behave like this, and its every exactly 1/4th of the time in reality

OpenStudy (dan815):

and its never* exactly 1/4th of the time in reality.

OpenStudy (dan815):

ok wait, i dont mean to confuse you, but it can be 1/4th of the time in experimental probability, however, its very unlikely its 1/4th of the time always. and you need to perform the experiment an infinite number of times to show that it is really going to 1/4th, and we cannot do an infinite experiments in reality. So this has to be theoretical

OpenStudy (anonymous):

I got it, hahaha thank you